among N2O,SO2,I2+ And I3-,the linear species are N2O and I3- ... why?

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

1.NO3-

x = 5 + 0 + 3
x = 8

Hybridisation is sp3.
So, geometry of the molecule is tetrahedral & shape is linear.

2.NO.

x = 5 + 0 + 0.

As x is between 4 & 6 hybridisation of the molecule is somewhere between sp & sp2.
The shape is bent.

3.NO2

x = 5 + 0 + 0.

Hybridisation & shape will be same as above.

4.N2

x = 5 + 1 + 0
x = 6

Hybridisation is sp2.
& geometry is trigonal planar whereas shape is linear.